OpenStudy (anonymous):

Solve the equation 8 = (4x + 8). Using complete sentences, explain the strategy used in solving this equation. (6 points)

OpenStudy (anonymous):

\[4x+8=8\] subtract 8 from both sides

OpenStudy (anonymous):

we get ???

OpenStudy (anonymous):

i am doing equational functions.

OpenStudy (anonymous):

do you know how to do those?

OpenStudy (anonymous):

sorry eqaution functions

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

lets say you have to write 5x+6y=-12 in function form. when you are writing an equation in function form, you usually solve for f(x) or g(x), but in this case, y is the same as f(x) or g(x). so to write 5x+6y=-12 in function form, you do this: 5x+6y=-12 6y= -5x -12 subtract 5x from both sides y= -5/6x -2 divide both sides by 6 i got it on this website http://wiki.answers.com/Q/How_do_you_rewrite_an_equation_in_function_form#ixzz1tji0vdRV

OpenStudy (anonymous):

ok thanks:D ttyl